### v1

@@ 3 The Algorithm
We use $W$ and $\mathbf{u}$ to denote the differentially privately released approximations of the sum of edge weights and the upper bound of each edge weight of $\hat{G}$ respectively. Then we reweigh the graph $\hat{G}$ to obtain $\overline{G}$ with adjacency matrix $\overline{\mathbf{A}}$, where the sum of edge weights is $W$. Specifically, we require that $W\geq\hat{W}$ and $\mathbf{u}\geq\overline{\mathbf{w}}$ with high probability. In the non-degenerate case when $\hat{W}$ and $\ell_{3}(\overline{G})$ are moderately large, we can guarantee that $W=\Theta(\hat{W})$ and ~$\tilde{\ell}_{3}(\hat{G})=\Theta(\ell_{3}(\hat{G}))=\Theta(\ell_{3}(\overline{G}))$. In the following, we treat $\overline{G}$ as the input graph with public $W$ and $\mathbf{u}$. More details are deferred to Appendix B.
Recall that $\mathbf{A}_{\triangle}$ is the adjacency matrix of the triangle-motif weighted graph of $G$. We let $\mathbf{D}^{(e)(t)}_{\triangle}$ denote the derivative of $\mathbf{A}_{\triangle}^{(t)}$ at $\mathbf{w}_{e}^{(t)}$, which is defined in Fact 15. For more details, refer to Section 2.2.
@@
Our algorithm will be based on the following optimization problem, which in turns is based on Section 4.1 that relates the triangle-motif cut difference and the following SDP form approximation. {mdframed}
$F_{\triangle}(\mathbf{w},\mathbf{X})$ $=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}\\ \mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}&\mathbf{0}\end{array}\right)\bullet\mathbf{X}+\lambda\log\det\mathbf{X}$ (5)
$+\sum_{(i,j)\in\binom{V}{2}}3(\mathbf{w}_{(i,j)}-\mathbf{\overline{w}}_{(i,j)})^{2}\cdot\sum_{s\in V\setminus\{i,j\}}(\mathbf{u}_{(i,s)}+\mathbf{u}_{(j,s)})$ (6)
$f_{\triangle}(\mathbf{w})=\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X}),$ (7)
where the domains of $\mathbf{w}$ and $\mathbf{X}$ are defined as follows:
$\mathcal{X}=\{\mathbf{w}\in\mathbb{R}_{+}^{\binom{V}{2}}:\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}=W,\mathbf{w}_{e}\leq\mathbf{u}_{e}\ \mathrm{for}\ \forall e\}$ (8)
$\mathcal{D}=\left\{\mathbf{X}\in\mathbb{R}^{2n}:\mathbf{X}\ \mathrm{is}\ \mathrm{symmetric},\mathbf{X}\succeq\frac{1}{n}\mathbf{I}_{2n},\mathrm{and}\ \mathbf{X}_{ii}=1\ \mathrm{for}\ \forall{i}\right\}$ (9)
Note that, the independent variable of our function is $\mathbf{w}$, while $\mathbf{A}_{\triangle}$ is used in Equation 5. However, $(\begin{array}[]{cc}\mathbf{0}&\mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}\\ \mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}&\mathbf{0}\end{array})\bullet\mathbf{X}$ is neither private nor convex. Therefore, we use the regularizer $\lambda\log\det\mathbf{X}$ to control the stability and $\sum_{(i,j)\in\binom{V}{2}}3(\mathbf{w}_{(i,j)}-\mathbf{\overline{w}}_{(i,j)})^{2}\cdot\sum_{s\in V\setminus\{i,j\}}(\mathbf{u}_{(i,s)}+\mathbf{u}_{(j,s)})$ to control the convexity.
@@
Now, we state our algorithm (Algorithm 2), which iteratively and privately solves the above optimization problem and outputs a graph $G$ that approximates each triangle-motif cut of $\hat{G}$. Specifically, our algorithm is an instantiation of the stochastic mirror descent algorithm (Algorithm 1), that provides an iterative approach to solve an optimization problem defined by a convex function $f$. In each iteration, the algorithm computes an unbiased estimator $g$ of the gradient $\nabla f$, and then updates the solution based on $g$ and the mirror function $\Phi$.
@@
Denote $N(\mathbf{x},\mathbf{\Sigma})$ as the multivariate normal distribution with mean $\mathbf{x}$ and covariance matrix $\Sigma$. Additionally, ~$\tilde{\ell}_{3}(\hat{G})$, $U_{\triangle}$, $U_{\Lambda}$ are some quantities defined in Appendix B.
@@
11:  Set $T=\Theta(\frac{W\cdot(\varepsilon U_{\triangle}+U_{\Lambda})}{n\log(n/\delta)\tilde{\ell}_{3}(\hat{G})})$,$L=\log_{3}(\frac{3}{\beta})$, $\lambda=\Theta(\varepsilon^{-1})\tilde{\ell}_{3}(\hat{G})\sqrt{T}\log^{\frac{3}{2}}(\frac{T}{\delta})\log(\frac{3}{\beta})$, $\varepsilon_{1}=\varepsilon_{2}=\varepsilon_{3}=\frac{\varepsilon}{6}$, $\varepsilon_{4}=\frac{\varepsilon}{6L}$
@@
46:        Find the maximizer $\mathbf{X}^{(t)}=\operatorname*{arg\,max}_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w}^{(t)},\mathbf{X})$, where $F_{\triangle}$ is defined in Equation 5
@@
In Line 9 of Algorithm 2, we use the algorithm $\mathrm{MD\_Update}$ to update our solution. We now describe this mirror descent update step. Given a weight vector $\mathbf{w}^{(t)}$ for some $t\geq 1$, the estimated gradient $\mathbf{g}^{(t)}$, and $W,\mathbf{u},\eta$, we update $\mathbf{w}^{(t)}$ to $\mathbf{w}^{(t+1)}$ using Algorithm 3.
@@
The above algorithm corresponds to the stochastic mirror descent framework from Algorithm 1 by setting $f(x)=f_{\triangle}(\mathbf{w})$ which is defined in Equation 7, and $\Phi(x)=\Phi(\mathbf{w})=\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}\log(\mathbf{w}_{e})$.
@@ 4 Analysis of the Algorithm
In this section, we provide the analysis of our algorithm (Algorithm 2) and prove the correctness of Algorithm 2. In Section 4.1, we derive the SDP approximation of the triangle-motif cut difference. In Section 4.2, we derive the gradient of $f_{\triangle}$ and prove its convexity. Then in Section 4.3, we prove the correctness of Algorithm 3 for mirror descent update step. Then we give privacy analysis in Section 4.4 and determines the value of $\lambda$ to guarantee the $(\varepsilon,\delta)$-DP property. Section 4.5 contains the utility analysis, in which we show the difference between $f_{\triangle}$ and cut norm, and use Corollary 20 to determine the value of $T$ and bound the additive error. Finally, in Section 4.6, we analyze the running time of the algorithm.
@@ 4.1 The SDP Approximation
Inspired by the setting of edge cut EKKL (20) described in Section 2.3, we have the following method to bound the additive error for triangle-motif cut. Consider graphs $G_{1}$ and $G_{2}$. Let $(\mathbf{A}_{1})_{\triangle}$ and $(\mathbf{A}_{2})_{\triangle}$ be their triangle adjacency matrices. Then for a fixed cut $(S,V\setminus S)$, the $\triangle$-motif size of $G_{1}$ is $\sum_{v\in S,u\in V\setminus S}\left((\mathbf{A}_{1})_{\triangle}\right)_{u,v}$ and the $\triangle$-motif size of $G_{2}$ is $\sum_{v\in S,u\in V\setminus S}\left((\mathbf{A}_{2})_{\triangle}\right)_{u,v}$. So we can see that the triangle-motif cut difference between $G_{1}$ and $G_{2}$ is twice of the following expression:
@@
According to FK (99), the cut on a graph can be bounded by the cut norm which is defined in Section 2.3. Thus by an approximation in AN (04) and Lemma 18, we have the following lemma:
@@
To sum up, the triangle-motif cut difference between $G_{1}$ and $G_{2}$ can be bounded by Equation 14 up to a constant factor.
@@ 4.2 Gradient and Convexity
Recall $\mathcal{D}$ is the domain of $\mathbf{X}$, which is defined in Equation 9. We first state some of its useful properties.
@@
=$=\begin{cases}6\sum_{s\in V\setminus\{i,j\}}(\mathbf{u}_{(i,s)}+\mathbf{u}_{(j,s)}),&(i,j)=(k,\ell)\\ \mathbf{w}_{(j,\ell)}(\sum_{i^{\prime},j^{\prime}\in\{i,j,\ell\}}\mathbf{X}_{i^{\prime},n+j^{\prime}}+\sum_{i^{\prime},j^{\prime}\in\{i,j,\ell\}}\mathbf{X}_{n+i^{\prime},j^{\prime}}),&i=k\ \mathrm{and}\ j\neq\ell\\ 0,&o.w.\end{cases}$
@@
Thus, by Fact 28, for fixed $\mathbf{X}\in\mathcal{D}$, $\nabla_{\mathbf{w}}^{2}F_{\triangle}(\mathbf{w},\mathbf{X})$ is semi-definite, so $F_{\triangle}(\mathbf{w},\mathbf{X})$ is convex. By Theorem 27, $f_{\triangle}(\mathbf{w})=\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X})$ is convex with respect to $\mathbf{w}$.
Moreover, for any fixed $\mathbf{X}\in\mathcal{D}$, it can be verified that $F_{\triangle}(\mathbf{w},\mathbf{X})$ is differentiable at $\mathbf{w}$. Therefore, by Theorem 27, $f_{\triangle}(\mathbf{w})$ is differentiable. Furthermore,
@@ 4.3 Mirror Descent Update
We now show that the update method Algorithm 3 correctly implements the mirror descent framework in our setting. To do so, we note that the mirror function in our algorithm is $\Phi(\mathbf{w})=\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}\log(\mathbf{w}_{e})$, and thus it suffices to prove that the output of Algorithm 3 is indeed the minimizer of $D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$, where $\mathbf{y}^{(t+1)}$ is the vector satisfying the Line 1 in Algorithm 3, i.e., $\mathbf{y}^{(t+1)}_{e}=\mathbf{w}^{(t)}_{e}\exp(-\eta\mathbf{g}^{(t)}_{e})$ for $\forall e\in\binom{V}{2}$.
@@
Denote the output of Algorithm 3 as $\mathbf{w}^{(t+1)}$. Then it holds that $D_{\Phi}(\mathbf{w}^{(t+1)},\mathbf{y}^{(t+1)})=\min_{\mathbf{w}\in\mathcal{X}}D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$, where
@@
Recall $\mathcal{X}=\{\mathbf{w}\in\mathbb{R}_{+}^{\binom{V}{2}}:\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}=W,\mathbf{w}_{e}\leq\mathbf{u}_{e}\ \mathrm{for}\ \forall e\}$. Because $\mathbf{u}_{e}=\overline{\mathbf{w}}_{e}+\mathrm{Lap}(3/\varepsilon)+3\log(3n^{2}/\beta)/\varepsilon+\frac{W}{\binom{n}{2}}>\frac{W}{\binom{n}{2}}$ for any $e\in\binom{V}{2}$ with high probability, $\mathbf{w}$ is an inner point of $\mathcal{X}$ when $\mathbf{w}_{e}=\frac{W}{\binom{n}{2}},\forall e\in\binom{V}{2}$, i.e., making the inequality constraints in $\mathcal{X}$ strictly feasible. It implies that the optimization satisfies Slater’s condition. Besides, it can be verified that $D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$ is a convex function with respect to $w$ Thus by Lemma 21, the following conditions are the necessary and sufficient condition for the optimal solution:
@@
where Equation (21) is equivalent to $01+\log(\mathbf{w}_{e})-1-\log(\mathbf{y}_{e}^{(t+1)})+\lambda_{e}+\mu=$ 0. Therefore we have,
@@
Denote $S=\sum_{e\in\binom{V}{2}}\mathbf{y}_{e}^{(t+1)}$. Assume $\mathbf{w}^{*}\in\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{X}}\min_{\mathbf{w}\in\mathcal{X}}D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$ is a solution with $k$ coordinates =$\mathcal{I}=\{i_{1},\dots,i_{k}\}$ satisfying the equality condition of Equation (24), i.e., $\mathbf{w}_{e_{i}}=\mathbf{u}_{e_{i}}$, for any $i\in\mathcal{I}$. We will consider two cases: $k=0$ and $k\neq 0$.
@@
When $k=0$, we have $\lambda_{e}=0$ for $\forall e\in\binom{V}{2}$. Thus, by Equation 26,
@@
Since there is no coordinate satisfying the equality condition of Equation (24), we have $\mathbf{w}^{*}_{e}<\mathbf{u}_{e}$ for any $e$. Let $e_{1},e_{2},\dots,e_{N}$ be the ordered sequence of edges given in Algorithm 3. Note that by the above discussion, $W_{1}=W$, $S_{1}=S$, and
@@
Then by induction we have that $W_{i+1}=W_{1}-\sum_{\ell=1}^{i}\mathbf{w}_{e_{\ell}}^{(t+1)}$ and $S_{i+1}=S_{1}-\sum_{\ell=1}^{i}\mathbf{y}_{e_{\ell}}^{(t+1)}$, which further implies that $W_{i+1}=S_{i+1}=\exp(-\mu)$. Thus, $\mathbf{w}_{e_{i+1}}^{(t+1)}=\frac{W_{i+1}\cdot\mathbf{y}_{e_{{i}}}^{(t+1)}}{S_{i+1}}=\exp(-\mu)\cdot\mathbf{y}_{e_{{i}}}^{(t+1)}=\mathbf{w}_{e_{i+1}}^{*}$. Therefore, the output solution $\mathbf{w}^{(t+1)}$ of Algorithm 3 is indeed the optimum solution $\mathbf{w}^{*}$.
Case 2: In the following we consider the case that $k\neq 0$. In this case, let =$\mathcal{I}=\{i_{1},\dots,i_{k}\}$ denote the index set of the $k$ coordinates such that $\mathbf{w}^{*}_{e_{i}}=\mathbf{u}_{e_{i}}$, for any $i\in\mathcal{I}$, and let $\mathcal{J}=\{j_{1},\dots,j_{\binom{n}{2}-k}\}$ deonte the set of remaining indices.
@@
Then by Equation (23), Equation (24) and Equation (26), we have $\mathbf{y}_{e_{j}}^{(t+1)}e^{-\mu}<\mathbf{u}_{e_{j}}$, for any $j\in\mathcal{J}$, and $e^{-\mu}=\frac{W-\sum_{i\in\mathcal{I}}\mathbf{u}_{e_{i}}}{S-\sum_{i\in\mathcal{I}}\mathbf{y}_{e_{i}}^{(t+1)}}$. Then by Equation (25) and Equation (26), we have $\mathbf{u}_{e_{i}}=\mathbf{w}^{*}_{e_{i}}\leq\mathbf{y}_{e_{i}}^{(t+1)}e^{-\mu}$ for any $i\in\mathcal{I}$. That is, $\frac{\mathbf{y}_{e_{j}}^{(t+1)}}{\mathbf{u}_{e_{j}}}<e^{\mu}\leq\frac{\mathbf{y}_{e_{i}}^{(t+1)}}{\mathbf{u}_{e_{i}}}$, for any $i\in\mathcal{I}$ and $j\in\mathcal{J}$. Thus, all the edges with indices in $\mathcal{I}$ appear before those with indices in in $\mathcal{J}$ according to the non-decreasing order in Algorithm 3. That is,
=$\mathcal{I}=\{1,\dots,k\},\mathcal{J}=\{k+1,\dots,N\},\text{ and $\frac{\mathbf{y}^{(t+1)}_{e_{1}}}{\mathbf{u}_{e_{1}}}\geq\dots\geq\frac{\mathbf{y}^{(t+1)}_{e_{k}}}{\mathbf{u}_{e_{k}}}\geq\exp(\mu)>\frac{\mathbf{y}^{(t+1)}_{e_{k+1}}}{\mathbf{u}_{e_{k+1}}}\dots\geq\frac{\mathbf{y}^{(t+1)}_{e_{N}}}{\mathbf{u}_{e_{N}}}$.}$
Then by Algorithm 3, $W_{1}=W$, and $S_{1}=S$. Furthermore, for each $1\leq\ell\leq i\leq k$, $\exp(-\mu)\geq\frac{\mathbf{u}_{e_{i}}}{\mathbf{y}_{e_{i}}^{(t+1)}}\geq\frac{\mathbf{u}_{e_{\ell}}}{\mathbf{y}_{e_{\ell}}^{(t+1)}}$. Thus, by the fact that $\frac{a_{1}+b_{1}}{a_{2}+b_{2}}\geq\min\{\frac{a_{1}}{a_{2}},\frac{b_{1}}{b_{2}}\}$ for any $a_{1},b_{1}\geq 0,a_{2},b_{2}>0$, we have that

### v2

@@ 3 The Algorithm
[…]
Recall that $\mathbf{A}_{\triangle}$ is the adjacency matrix of the triangle-motif weighted graph of $G$. We let $\mathbf{D}^{(e)(t)}_{\triangle}$ denote the derivative of $\mathbf{A}_{\triangle}^{(t)}$ at $\mathbf{w}_{e}^{(t)}$, which is defined in Fact 15. For more details, refer to Section˜2.2.
@@
Our algorithm will be based on the following optimization problem, which in turns is based on Section˜4.1 that relates the triangle-motif cut difference and the following SDP form approximation. $F_{\triangle}(\mathbf{w},\mathbf{X})$ $=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}\\ \mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}&\mathbf{0}\end{array}\right)\bullet\mathbf{X}+\lambda\log\det\mathbf{X}$ (5) $+\sum_{(i,j)\in\binom{V}{2}}3(\mathbf{w}_{(i,j)}-\mathbf{\overline{w}}_{(i,j)})^{2}\cdot\sum_{s\in V\setminus\{i,j\}}(\mathbf{u}_{(i,s)}+\mathbf{u}_{(j,s)})$ (6) $f_{\triangle}(\mathbf{w})=\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X}),$ (7) where the domains of $\mathbf{w}$ and $\mathbf{X}$ are defined as follows: $\mathcal{X}=\{\mathbf{w}\in\mathbb{R}_{+}^{\binom{V}{2}}:\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}=W,\mathbf{w}_{e}\leq\mathbf{u}_{e}\ \mathrm{for}\ \forall e\}$ (8) $\mathcal{D}=\left\{\mathbf{X}\in\mathbb{R}^{2n}:\mathbf{X}\ \mathrm{is}\ \mathrm{symmetric},\mathbf{X}\succeq\frac{1}{n}\mathbf{I}_{2n},\mathrm{and}\ \mathbf{X}_{ii}=1\ \mathrm{for}\ \forall{i}\right\}$ (9)
Note that, the independent variable of our function is $\mathbf{w}$, while $\mathbf{A}_{\triangle}$ is used in Equation˜5. However, $(\begin{array}[]{cc}\mathbf{0}&\mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}\\ \mathbf{A}_{\triangle}-\mathbf{\overline{A}}_{\triangle}&\mathbf{0}\end{array})\bullet\mathbf{X}$ is neither private nor convex. Therefore, we use the regularizer $\lambda\log\det\mathbf{X}$ to control the stability and $\sum_{(i,j)\in\binom{V}{2}}3(\mathbf{w}_{(i,j)}-\mathbf{\overline{w}}_{(i,j)})^{2}\cdot\sum_{s\in V\setminus\{i,j\}}(\mathbf{u}_{(i,s)}+\mathbf{u}_{(j,s)})$ to control the convexity.
@@
Now, we state our algorithm (Algorithm˜2), which iteratively and privately solves the above optimization problem and outputs a graph $G$ that approximates each triangle-motif cut of $\hat{G}$. Specifically, our algorithm is an instantiation of the stochastic mirror descent algorithm (Algorithm˜1), that provides an iterative approach to solve an optimization problem defined by a convex function $f$. In each iteration, the algorithm computes an unbiased estimator $g$ of the gradient $\nabla f$, and then updates the solution based on $g$ and the mirror function $\Phi$.
We invoke Algorithm˜1 with $f(x)=f_{\triangle}(\mathbf{w})$, $\Phi(x)=\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}\log(\mathbf{w}_{e})$, and mirror update step from Algorithm˜3. Additionally, we set $\left\|\cdot\right\|$ as $l_{1}$ norm, hence $\left\|\cdot\right\|_{*}$ is $l_{\infty}$ norm.
@@
[…]
@@
46:   Find the maximizer $\mathbf{X}^{(t)}=\operatorname*{arg\,max}_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w}^{(t)},\mathbf{X})$, where $F_{\triangle}$ is defined in Equation˜5
@@
In Line 9 of Algorithm˜2, we use the algorithm $\mathrm{MD\_Update}$ to update our solution. We now describe this mirror descent update step. Given a weight vector $\mathbf{w}^{(t)}$ for some $t\geq 1$, the estimated gradient $\mathbf{g}^{(t)}$, and $W,\mathbf{u},\eta$, we update $\mathbf{w}^{(t)}$ to $\mathbf{w}^{(t+1)}$ using Algorithm˜3.
@@
The above algorithm corresponds to the stochastic mirror descent framework from Algorithm˜1 by setting $f(x)=f_{\triangle}(\mathbf{w})$ which is defined in Equation˜7, and $\Phi(x)=\Phi(\mathbf{w})=\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}\log(\mathbf{w}_{e})$.
@@ 4 Analysis of the Algorithm
In this section, we provide the analysis of our algorithm (Algorithm˜2) and prove the correctness of Algorithm˜2. In Section˜4.1, we derive the SDP approximation of the triangle-motif cut difference. In Section˜4.2, we derive the gradient of $f_{\triangle}$ and prove its convexity. Then in Section˜4.3, we prove the correctness of Algorithm˜3 for mirror descent update step. Then we give privacy analysis in Section˜4.4 and determines the value of $\lambda$ to guarantee the $(\varepsilon,\delta)$-DP property. Section˜4.5 contains the utility analysis, in which we show the difference between $f_{\triangle}$ and cut norm, and use Corollary 20 to determine the value of $T$ and bound the additive error. Finally, in Section˜4.6, we analyze the running time of the algorithm.
@@ 4.1 The SDP Approximation
Inspired by the setting of edge cut EKKL (20) described in Section˜2.3, we have the following method to bound the additive error for triangle-motif cut. Consider graphs $G_{1}$ and $G_{2}$. Let $(\mathbf{A}_{1})_{\triangle}$ and $(\mathbf{A}_{2})_{\triangle}$ be their triangle adjacency matrices. Then for a fixed cut $(S,V\setminus S)$, the $\triangle$-motif size of $G_{1}$ is $\sum_{v\in S,u\in V\setminus S}\left((\mathbf{A}_{1})_{\triangle}\right)_{u,v}$ and the $\triangle$-motif size of $G_{2}$ is $\sum_{v\in S,u\in V\setminus S}\left((\mathbf{A}_{2})_{\triangle}\right)_{u,v}$. So we can see that the triangle-motif cut difference between $G_{1}$ and $G_{2}$ is twice of the following expression:
@@
According to FK (99), the cut on a graph can be bounded by the cut norm which is defined in Section˜2.3. Thus by an approximation in AN (04) and Lemma 18, we have the following lemma:
@@
To sum up, the triangle-motif cut difference between $G_{1}$ and $G_{2}$ can be bounded by Equation˜10 up to a constant factor.
@@ 4.2 Gradient and Convexity
Recall $\mathcal{D}$ is the domain of $\mathbf{X}$, which is defined in Equation˜9. We first state some of its useful properties.
@@
$=\begin{cases}6\sum_{s\in V\setminus\{i,j\}}(\mathbf{u}_{(i,s)}+\mathbf{u}_{(j,s)}),&(i,j)=(k,\ell)\\ \mathbf{w}_{(j,\ell)}(\sum_{i^{\prime},j^{\prime}\in\{i,j,\ell\}}\mathbf{X}_{i^{\prime},n+j^{\prime}}+\sum_{i^{\prime},j^{\prime}\in\{i,j,\ell\}}\mathbf{X}_{n+i^{\prime},j^{\prime}}),&i=k\ \mathrm{and}\ j\neq\ell\\ 0,&o.w.\end{cases}$
@@
Thus, by Fact 28, for fixed $\mathbf{X}\in\mathcal{D}$, $\nabla_{\mathbf{w}}^{2}F_{\triangle}(\mathbf{w},\mathbf{X})$ is semi-definite, so $F_{\triangle}(\mathbf{w},\mathbf{X})$ is convex. By Theorem˜27, $f_{\triangle}(\mathbf{w})=\max_{\mathbf{X}\in\mathcal{D}}F_{\triangle}(\mathbf{w},\mathbf{X})$ is convex with respect to $\mathbf{w}$.
Moreover, for any fixed $\mathbf{X}\in\mathcal{D}$, it can be verified that $F_{\triangle}(\mathbf{w},\mathbf{X})$ is differentiable at $\mathbf{w}$. Therefore, by Theorem˜27, $f_{\triangle}(\mathbf{w})$ is differentiable. Furthermore,
@@ 4.3 Mirror Descent Update
We now show that the update method Algorithm˜3 correctly implements the mirror descent framework in our setting. To do so, we note that the mirror function in our algorithm is $\Phi(\mathbf{w})=\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}\log(\mathbf{w}_{e})$, and thus it suffices to prove that the output of Algorithm˜3 is indeed the minimizer of $D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$, where $\mathbf{y}^{(t+1)}$ is the vector satisfying the Line 1 in Algorithm˜3, i.e., $\mathbf{y}^{(t+1)}_{e}=\mathbf{w}^{(t)}_{e}\exp(-\eta\mathbf{g}^{(t)}_{e})$ for $\forall e\in\binom{V}{2}$.
@@
Denote the output of Algorithm˜3 as $\mathbf{w}^{(t+1)}$. Then it holds that $D_{\Phi}(\mathbf{w}^{(t+1)},\mathbf{y}^{(t+1)})=\min_{\mathbf{w}\in\mathcal{X}}D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$, where
@@
Recall $\mathcal{X}=\{\mathbf{w}\in\mathbb{R}_{+}^{\binom{V}{2}}:\sum_{e\in\binom{V}{2}}\mathbf{w}_{e}=W,\mathbf{w}_{e}\leq\mathbf{u}_{e}\ \mathrm{for}\ \forall e\}$. Because $\mathbf{u}_{e}=\overline{\mathbf{w}}_{e}+\mathrm{Lap}(3/\varepsilon)+3\log(3n^{2}/\beta)/\varepsilon+\frac{W}{\binom{n}{2}}>\frac{W}{\binom{n}{2}}$ for any $e\in\binom{V}{2}$ with high probability, $\mathbf{w}$ is an inner point of $\mathcal{X}$ when $\mathbf{w}_{e}=\frac{W}{\binom{n}{2}},\forall e\in\binom{V}{2}$, i.e., making the inequality constraints in $\mathcal{X}$ strictly feasible. It implies that the optimization satisfies Slater’s condition. Besides, it can be verified that $D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$ is a convex function with respect to $w$ . Thus by Lemma 21, the following conditions are the necessary and sufficient condition for the optimal solution:
@@
where Equation (13) is equivalent to $1+\log(\mathbf{w}_{e})-1-\log(\mathbf{y}_{e}^{(t+1)})+\lambda_{e}+\mu=0$. Therefore we have,
@@
Denote $S=\sum_{e\in\binom{V}{2}}\mathbf{y}_{e}^{(t+1)}$. Assume $\mathbf{w}^{*}\in\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{X}}\min_{\mathbf{w}\in\mathcal{X}}D_{\Phi}(\mathbf{w},\mathbf{y}^{(t+1)})$ is a solution with $k$ coordinates $\mathcal{I}=\{i_{1},\dots,i_{k}\}$ satisfying the equality condition of Equation (16), i.e., $\mathbf{w}_{e_{i}}=\mathbf{u}_{e_{i}}$, for any $i\in\mathcal{I}$. We will consider two cases: $k=0$ and $k\neq 0$.
@@
When $k=0$, we have $\lambda_{e}=0$ for $\forall e\in\binom{V}{2}$. Thus, by Equation˜18,
@@
Since there is no coordinate satisfying the equality condition of Equation (16), we have $\mathbf{w}^{*}_{e}<\mathbf{u}_{e}$ for any $e$. Let $e_{1},e_{2},\dots,e_{N}$ be the ordered sequence of edges given in Algorithm˜3. Note that by the above discussion, $W_{1}=W$, $S_{1}=S$, and
@@
Then by induction we have that $W_{i+1}=W_{1}-\sum_{\ell=1}^{i}\mathbf{w}_{e_{\ell}}^{(t+1)}$ and $S_{i+1}=S_{1}-\sum_{\ell=1}^{i}\mathbf{y}_{e_{\ell}}^{(t+1)}$, which further implies that $W_{i+1}=S_{i+1}=\exp(-\mu)$. Thus, $\mathbf{w}_{e_{i+1}}^{(t+1)}=\frac{W_{i+1}\cdot\mathbf{y}_{e_{{i}}}^{(t+1)}}{S_{i+1}}=\exp(-\mu)\cdot\mathbf{y}_{e_{{i}}}^{(t+1)}=\mathbf{w}_{e_{i+1}}^{*}$. Therefore, the output solution $\mathbf{w}^{(t+1)}$ of Algorithm˜3 is indeed the optimum solution $\mathbf{w}^{*}$.
Case 2: In the following we consider the case that $k\neq 0$. In this case, let $\mathcal{I}=\{i_{1},\dots,i_{k}\}$ denote the index set of the $k$ coordinates such that $\mathbf{w}^{*}_{e_{i}}=\mathbf{u}_{e_{i}}$, for any $i\in\mathcal{I}$, and let $\mathcal{J}=\{j_{1},\dots,j_{\binom{n}{2}-k}\}$ deonte the set of remaining indices.
@@
Then by Equation (15), Equation (16) and Equation (18), we have $\mathbf{y}_{e_{j}}^{(t+1)}e^{-\mu}<\mathbf{u}_{e_{j}}$, for any $j\in\mathcal{J}$, and $e^{-\mu}=\frac{W-\sum_{i\in\mathcal{I}}\mathbf{u}_{e_{i}}}{S-\sum_{i\in\mathcal{I}}\mathbf{y}_{e_{i}}^{(t+1)}}$. Then by Equation (17) and Equation (18), we have $\mathbf{u}_{e_{i}}=\mathbf{w}^{*}_{e_{i}}\leq\mathbf{y}_{e_{i}}^{(t+1)}e^{-\mu}$ for any $i\in\mathcal{I}$. That is, $\frac{\mathbf{y}_{e_{j}}^{(t+1)}}{\mathbf{u}_{e_{j}}}<e^{\mu}\leq\frac{\mathbf{y}_{e_{i}}^{(t+1)}}{\mathbf{u}_{e_{i}}}$, for any $i\in\mathcal{I}$ and $j\in\mathcal{J}$. Thus, all the edges with indices in $\mathcal{I}$ appear before those with indices in in $\mathcal{J}$ according to the non-decreasing order in Algorithm˜3. That is,
$\mathcal{I}=\{1,\dots,k\},\mathcal{J}=\{k+1,\dots,N\},\text{ and $\frac{\mathbf{y}^{(t+1)}_{e_{1}}}{\mathbf{u}_{e_{1}}}\geq\dots\geq\frac{\mathbf{y}^{(t+1)}_{e_{k}}}{\mathbf{u}_{e_{k}}}\geq\exp(\mu)>\frac{\mathbf{y}^{(t+1)}_{e_{k+1}}}{\mathbf{u}_{e_{k+1}}}\dots\geq\frac{\mathbf{y}^{(t+1)}_{e_{N}}}{\mathbf{u}_{e_{N}}}$.}$
Then by Algorithm˜3, $W_{1}=W$, and $S_{1}=S$. Furthermore, for each $1\leq\ell\leq i\leq k$, $\exp(-\mu)\geq\frac{\mathbf{u}_{e_{i}}}{\mathbf{y}_{e_{i}}^{(t+1)}}\geq\frac{\mathbf{u}_{e_{\ell}}}{\mathbf{y}_{e_{\ell}}^{(t+1)}}$. Thus, by the fact that $\frac{a_{1}+b_{1}}{a_{2}+b_{2}}\geq\min\{\frac{a_{1}}{a_{2}},\frac{b_{1}}{b_{2}}\}$ for any $a_{1},b_{1}\geq 0,a_{2},b_{2}>0$, we have that